OpenStudy (anonymous):

van is 75% than kieth .if van's heightt increasess by 40% while kieth's increases only by 25%, by what percent is van taller than kieth now? 96%, 70.00%, 81%,62.00%

OpenStudy (anonymous):

here is an easy trick to use for problems like this. since you are not told what either height is, lets put Keith's height at 100 and 75% of 100 is 75, so Van's height is 175

OpenStudy (anonymous):

that was wrong. Van increased by 40% and \[1.4\times 175=254\] whereas Keith increased by 25% and \[1.25\times 100=125\]

OpenStudy (anonymous):

so either i read the problem wrong or there is something missing.

OpenStudy (anonymous):

ok let me check

OpenStudy (anonymous):

what was this line van is 75% than kieth ??

OpenStudy (anonymous):

opps sorry buddy right is this van is 75% taller than keith. if van's height increases by 40% while kieth's increases only by 25%, by what percent is van taller than kieth now?

OpenStudy (anonymous):

satellite dude help me

OpenStudy (anonymous):

ok lets start again

OpenStudy (anonymous):

van is 75% taller than keith. we put keith at 100, van at 175 keith increases by 25% he is 125 van increases by 40% he is 245

OpenStudy (anonymous):

245-125=120 and \[\frac{120}{125}=.96=96\%\]

OpenStudy (anonymous):

thnks dude genius answer